Today was rainy, with heavy falls early, gradually easing through the day to occasional scattered showers by mid afternoon. It was also Jen, Gary and Eckhard's last day, many thanks for their efforts. Gary got network access working at both sites (via cell modems); although they are not yet transferring data automatically, it is possible to remotely access the two sites. Eckhard did more checks on the sodar-RASS, it is working very well and seems to be unaffected by the nearby high tension power lines. The wind profiler is also working well with reduced clutter since installing the clutter screen. The heavy rain revealed a couple of minor problems, a small leak at the profiler shelter and an intermittent connection to the webcam, both of which will be repaired as conditions dry out.